Question 20. The answer is wrong

A rectangular coil with sides 0.130 m by 0.230 m h
Answer #1


area of the coil, A=0.13*0.23 m^2


no of turns, N=487


magnetic field, B=0.52T


emf=115 v


use,


emf=N*B*A*w


115=487*0.52*(0.13*0.23)*(2pi*f)


===> f=2.417 Hz


frequency f=2.42 Hz